OpenStudy (anonymous):

If x is a real number, which of the following statements is correct about the equation 12x+4/2= 24x+16/4?

OpenStudy (anonymous):

a) The equation is always true b)the equation is always true, except when x=-1/3 or x=-2/3 c)the equation is sometimes true d) the equation is never true

OpenStudy (anonymous):

\[12x+4/2=24x+16/4\] group like terms\[12x-24x=16/4-4/2\] \[-12x=2\] \[x=2/-12\] \[x=-0.167\]

OpenStudy (anonymous):

so can u guess it now

OpenStudy (anonymous):

@ludvic